# DogeRM: Equipping Reward Models with Domain Knowledge through Model Merging

### v1

@@ Abstract
Reinforcement learning from human feedback (RLHF) is a popular strategy for aligning large language models (LLMs) with desired behaviors. Reward modeling is a crucial step in RLHF. However, collecting paired preference data for training reward models is often costly and time-consuming, especially for domain-specific preferences requiring expert annotation. To address this challenge, we propose the Domain knowledge merged Reward Model (DogeRM), a novel framework that integrates domain-specific knowledge into a general reward model by model merging. The experiments demonstrate that DogeRM enhances performance across different benchmarks and provide a detailed analysis showcasing the effects of model merging, showing the great potential of facilitating model alignment.
@@
Tzu-Han Lin  Chen-An Li  Hung-yi Lee  Yun-Nung Chen National Taiwan University, Taipei, Taiwan {r12944034,b08902123}@csie.ntu.edu.twhungyilee@ntu.edu.twy.v.chen@ieee.org
@@ 1 Introduction
Modern large language models (LLMs), such as GPT-4 Achiam et al. (2023) and Gemini Team et al. (2023), showcase impressive capabilities across various tasks Gao et al. (2023); Beeching et al. (2023), with aligning their behavior with human preferences. Reinforcement learning from human feedback (RLHF) is a prominent technique for enhancing the alignment of desired behaviors in LLMs Christiano et al. (2017); Ziegler et al. (2019); Ouyang et al. (2022). A key component of RLHF is its reward models (RMs), which assess entire sentences generated by policy models. The reward signals produced by these RMs are instrumental in adjusting the parameters of the policy models, thus directly impacting the policy models’ effectiveness.
RMs are developed by training LLMs on paired preference data to simulate human judgment Ouyang et al. (2022). This preference data consists of two responses to a given user input, accompanied by a human-assigned label indicating which response is more preferred. However, gathering such preference data can be costly and time-consuming due to the requirement of human annotation Stiennon et al. (2020a). This challenge becomes more pronounced when handling domain-specific preference data, as it necessitates expertise from domain specialists.
Figure 1: The framework of DogeRM, where we merge the general RM with a domain-specific LM to create the domain-specific RM.
Recent developments have demonstrated the effectiveness of model merging techniques in strategically integrating multiple domain-specific models into a multi-domain model without requiring additional training Wortsman et al. (2022); Ilharco et al. (2023). Furthermore, domain-specific SFT data is relatively more accessible compared to preference data. Moreover, many high-quality domain-specific models are available on open-source platforms Wolf et al. (2019), which can be directly employed in the merging process. This brings us to consider a novel approach: Is it possible to equip reward models with domain knowledge through merging with domain-specific language models?
In this work, we propose Domain knowledge merged Reward Model (DogeRM), exploring the potential of merging a reward model trained on a general open-sourced preference dataset with a language model fine-tuned on domain-specific datasets, such as math and coding. An illustration of DogeRM is presented in Figure 1. We evaluate DogeRM using RewardBench Lambert et al. (2024), Auto-J Eval Li et al. (2024) and Best-of-N sampling on GSM8K Cobbe et al. (2021) and MBPP Austin et al. (2021). Our results demonstrate that DogeRM improves performance and can be generalized to different model architectures. We also conduct a comprehensive analysis to demonstrate the impact of model merging.
@@ 2 Related Work
RMs are crucial for aligning language models with human preferences, providing proxy rewards as training signals for policy models. Previous work has employed RL algorithms with RMs to guide language models towards human preferences in various NLP tasks Ziegler et al. (2020); Stiennon et al. (2020b); Wu et al. (2021); Nakano et al. (2022); Menick et al. (2022) and instruction-following Ouyang et al. (2022); Ramamurthy et al. (2023). In RLHF literature, RMs evaluate the quality of instructions and responses based on criteria like helpfulness and harmlessness Bai et al. (2022) or more fine-grained objectives Wu et al. (2023).
Several open-source paired preference datasets are available for training RMs for RLHF, such as OpenAI Summarization Stiennon et al. (2020b), HH-RLHF Bai et al. (2022), SHP Ethayarajh et al. (2022), Ultrafeedback Cui et al. (2023), PKU-SafeRLHF Ji et al. (2023), HelpSteer Wang et al. (2023), Nectar Zhu et al. (2023), and UltraInteract Yuan et al. (2024). However, most datasets are not domain-specific. To address this, our work focuses on merging RMs with domain-specific language models, aiming to equip RMs with domain knowledge.
@@
Model merging integrates multiple task-specific models into a single unified model without additional training. A straightforward approach involves averaging parameters from models fine-tuned from the same initial model Wortsman et al. (2022). Another method employs weighted averaging of model parameters Matena and Raffel (2022); Jin et al. (2023).
Another innovative approach involves creating task vectors by subtracting the weights of a pre-trained model from those of the same model after fine-tuning for a specific task. This method showcases the flexibility and composability of these vectors through arithmetic operations Ilharco et al. (2023); Yadav et al. (2024); Huang et al. (2024).
Some recent work focused on model merging to align with user preferences. They interpolated model parameters fine-tuned on diverse rewards Rame et al. (2024a); Jang et al. (2023); Wang et al. (2024a), or merging RMs for combining different aspects of rewards Rame et al. (2024b).
However, these approaches still require domain-specific preference data to integrate domain knowledge. Our work eliminates the requirement of domain-specific preference data and focuses on incorporating domain-specific knowledge into RMs through model merging.
@@ 3.1 Reward Modeling
$\mathcal{L}_{\text{RM}}=-\log{[\sigma(r(x,y_{c}))-\sigma(r(x,y_{r}))]}\vspace{-5pt}$ (1)
@@
Model Reward Bench Auto-J Eval
Chat Chat-Hard Safety Reasoning Code Math Others
@@
LLaMA-2 RM 95.8 47.6 44.6 78.9 68.2 76.2 84.4 79.2
Ours (+ MetaMath) 95.8 44.5 43.5 85.7 79.6 79.8 87.5 79.3
Ours (+ MAmmoTH) 96.1 44.7 43.8 84.1 85.2 79.8 87.5 79.7
Ours (+ Code Model) 96.1 45.6 43.9 84.3 71.8 82.1 87.5 79.7
Table 1: Performance comparison on various benchmarks.
@@ 3.2 Model Merging
$\theta^{\text{SFT}}=\{\theta^{\text{SFT}}_{\text{emb}},\theta^{\text{SFT}}_{\text{trans}},\theta^{\text{SFT}}_{\text{dec}}\}\vspace{-5pt}$ (2)
@@
$\theta^{\text{RM}}=\{\theta^{\text{RM}}_{\text{emb}},\theta^{\text{RM}}_{\text{trans}},\theta^{\text{RM}}_{\text{reg}}\}\vspace{-5pt}$ (3)
@@
where $t_{i}$ is a common token to both models, $\theta_{\text{emb},t_{i}}$ is the corresponding embedding, and $\lambda$ are weights for SFT parameters ranging from 0 to 1.
@@
$\theta^{\text{MERGE}}_{\text{emb},t_{i}}=\begin{cases}\theta^{\text{SFT}}_{\text{emb},t_{i}}&\text{If $t_{i}$ is unique to SFT model}\\ \theta^{\text{RM}}_{\text{emb},t_{i}}&\text{If $t_{i}$ is unique to RM}\end{cases}$ (5)
@@ 4.1 Experimental Setup
To fine-tune the backbone of our reward model, we utilize the 10k SFT split from Alpacafarm Dubois et al. (2023). For reward modeling, we employ the UltraFeedback Cui et al. (2023). The details of training and these datasets are presented in Appendix A and C, respectively.
@@
For the math LLMs, we adopt the open-source models, MetaMath-7B Yu et al. (2024) and MAmmoTH-7B Yue et al. (2024a), both of which are fine-tuned from LLaMA-2-7B. For code generation LLM, since we could not find open-source models with detailed training information, we use OSS-Instruct and Magicoder-Evol-Instruct Wei et al. (2024) to fine-tune LLaMA-2-7B ourselves. We refer to this model as the Code Model in the following sections. The details of code fine-tuning datasets and math models are presented in Appendix C and D, respectively.
@@ 4.2 Evaluation
We evaluate the reward models using two benchmarks, RewardBench Lambert et al. (2024) and Auto-J Eval Li et al. (2024). These benchmarks provide paired instruction-completion data, with the preferred completion annotated as chosen and the other as rejected, using accuracy as the evaluation metric. We use the core set of RewardBench, focusing primarily on the reasoning category to evaluate the model’s abilities in code and mathematical reasoning. For Auto-J Eval, we use pair-wise testing data and categorize the dataset into three categories: code, math, and others, following Yuan et al. 2024. Additionally, to further test our reward model’s effectiveness in enhancing model performance through reranking, we conduct best-of-N sampling on zero-shot prompted responses from llama-2-7B-chat for GSM8K Cobbe et al. (2021) and MBPP Austin et al. (2021). We present results using a weight factor of $\lambda=0.35$ for the main findings, with an analysis of the impact of $\lambda$ detailed in section 4.4 and results for different values of $\lambda$ presented in Appendix F.
@@ 4.3 Results
The main results RM benchmark are shown in Table 1. Merging the LLaMA-2 RM with MetaMath-7B Yu et al. (2024) and MAmmoTH-7B Yue et al. (2024a) improves math performance on RewardBench by 11.4% and 17%, respectively, and coding performance by 5.2% and 5.8%, respectively. Similar enhancements are seen on Auto-J Eval, with gains in both math and coding. Merging our LLaMA-2 RM with the Code Model further improves coding performance on RewardBench and Auto-J Eval by 5.4% and 6%, respectively, along with noticeable improvements in math performance on both benchmarks. Although our methods improve the reasoning domain, performance in other domains did not significantly degrade.
@@
Figure 2 shows the results, with accuracy improvements on GSM8K and MBPP. At the best-of-16 setting, merging with math models improves GSM8K by 5%, and merging with the Code Model enhances MBPP by 1.4%.
@@ 4.4 Analysis
To test the adaptability of DogeRM to different model architectures, we use an open-source Mistral-based Jiang et al. (2023) RM Ray2333 (2024) merging with Misrtral-based MAmmoTH2-7B-Plus Yue et al. (2024b). Details of these models are presented in Appendix D. The results for $\lambda=0.35$ in reasoning domains on RM benchmarks and best-of-N sampling on GSM8K, with N=16, are presented in Table 2. Our method improves math performance by 30% on RewardBench and 3% on Auto-J Eval. Additionally, we enhance reranking performance on GSM8K by 2.4%. These results demonstrate the adaptability of our methods to different model architectures. The results for different $\lambda$ are presented in Appendix F.
@@ 5 Conclusion
In this work, we introduce a novel approach, DogeRM, which integrates domain knowledge into RM by merging it with the SFT model. We demonstrate that DogeRM enhances performance on math and coding benchmarks and can be generalized to different architectures. A series of analyses show that DogeRM effectively affects the reward signal corresponding to chosen and rejected prompts.
@@
There are several limitations in our work: (1) We only tested our framework in the math and coding domains; other domains such as medicine, finance, and law have not been explored. (2) Our method was tested exclusively on 7B models, and we have not evaluated its performance on models of larger or smaller sizes. (3) While our framework is compatible with various merging techniques, such as TIES-Merge Yadav et al. (2024), we did not investigate the actual effects of these different methods.

### v2

@@ Abstract
Reinforcement learning from human feedback (RLHF) is a popular strategy for aligning large language models (LLMs) with desired behaviors. Reward modeling is a crucial step in RLHF. However, collecting paired preference data for training reward models is often costly and time-consuming, especially for domain-specific preferences requiring expert annotation. To address this challenge, we propose the Domain knowledge merged Reward Model (DogeRM), a novel framework that integrates domain-specific knowledge into a general reward model by model merging. The experiments demonstrate that DogeRM enhances performance across different benchmarks and provide a detailed analysis showcasing the effects of model merging, showing the great potential of facilitating model alignment. [^1] The source code and trained models are released at
https://github.com/MiuLab/DogeRM.
@@
Tzu-Han Lin  Chen-An Li  Hung-yi Lee  Yun-Nung Chen National Taiwan University, Taipei, Taiwan {r12944034,r13942069,hungyilee}@ntu.edu.twy.v.chen@ieee.org
@@ 1 Introduction
Modern large language models (LLMs), such as GPT-4 (Achiam et al., 2023) and Gemini (Team et al., 2023), showcase impressive capabilities across various tasks (Gao et al., 2023; Beeching et al., 2023), with aligning their behavior with human preferences. Reinforcement learning from human feedback (RLHF) is a prominent technique for enhancing the alignment of desired behaviors in LLMs (Christiano et al., 2017; Ziegler et al., 2020; Ouyang et al., 2022). A key component of RLHF is its reward models (RMs), which assess entire sentences generated by policy models. The reward signals produced by these RMs are instrumental in adjusting the parameters of the policy models, thus directly impacting the policy models’ effectiveness.
RMs are developed by training LLMs on paired preference data to simulate human judgment (Ouyang et al., 2022). This preference data consists of two responses to a given user input, accompanied by a human-assigned label indicating which response is more preferred. However, gathering such preference data can be costly and time-consuming due to the requirement of human annotation (Stiennon et al., 2020). This challenge becomes more pronounced when handling domain-specific preference data, as it necessitates expertise from domain specialists.
Figure 1: The framework of DogeRM, illustrating the merging of a general RM with a domain-specific LM to create a domain-specific RM. All icons used in this figure are sourced from https://www.flaticon.com/.
Recent developments have demonstrated the effectiveness of model merging techniques in strategically integrating multiple domain-specific models into a multi-domain model without requiring additional training (Wortsman et al., 2022; Ilharco et al., 2023). Furthermore, domain-specific SFT data is relatively more accessible compared to preference data. Moreover, many high-quality domain-specific models are available on open-source platforms (Wolf et al., 2020), which can be directly employed in the merging process. This brings us to consider a novel approach: Is it possible to equip reward models with domain knowledge through merging with domain-specific language models?
In this work, we propose Domain knowledge merged Reward Model (DogeRM), exploring the potential of merging a reward model trained on a general open-sourced preference dataset with a language model fine-tuned on domain-specific datasets, such as math and coding. An illustration of DogeRM is presented in Figure 1. We evaluate DogeRM using RewardBench (Lambert et al., 2024), Auto-J Eval (Li et al., 2024) and Best-of-N sampling on GSM8K (Cobbe et al., 2021) and MBPP (Austin et al., 2021). Our results demonstrate that DogeRM improves performance and can be generalized to different model architectures. We also conduct a comprehensive analysis to demonstrate the impact of model merging.
@@ 2 Related Work
RMs are crucial for aligning language models with human preferences, providing proxy rewards as training signals for policy models. Previous work has employed RL algorithms with RMs to guide language models towards human preferences in various NLP tasks (Ziegler et al., 2020; Stiennon et al., 2020; Wu et al., 2021; Nakano et al., 2022; Menick et al., 2022) and instruction-following (Ouyang et al., 2022; Ramamurthy et al., 2023). In RLHF literature, RMs evaluate the quality of instructions and responses based on criteria like helpfulness and harmlessness (Bai et al., 2022) or more fine-grained objectives (Wu et al., 2023).
Several open-source paired preference datasets are available for training RMs for RLHF, such as OpenAI Summarization (Stiennon et al., 2020), HH-RLHF (Bai et al., 2022), SHP (Ethayarajh et al., 2022), Ultrafeedback (Cui et al., 2024), PKU-SafeRLHF (Ji et al., 2023), HelpSteer (Wang et al., 2024c), Nectar (Zhu et al., 2023), and UltraInteract (Yuan et al., 2024). However, most datasets are not domain-specific. To address this, our work focuses on merging RMs with domain-specific language models, aiming to equip RMs with domain knowledge.
@@
Model merging integrates multiple task-specific models into a single unified model without additional training. A straightforward approach involves averaging parameters from models fine-tuned from the same initial model (Wortsman et al., 2022). Another method employs weighted averaging of model parameters (Matena and Raffel, 2022; Jin et al., 2023).
Another innovative approach involves creating task vectors by subtracting the weights of a pre-trained model from those of the same model after fine-tuning for a specific task. This method showcases the flexibility and composability of these vectors through arithmetic operations (Ilharco et al., 2023; Yadav et al., 2024; Huang et al., 2024).
Some recent work focused on model merging to align with user preferences. They interpolated model parameters fine-tuned on diverse rewards (Rame et al., 2024a; Jang et al., 2023; Wang et al., 2024a), or merging RMs for combining different aspects of rewards (Rame et al., 2024b).
However, these methods still rely heavily on substantial domain-specific preference data to integrate domain knowledge. In contrast, our approach significantly reduces the need for such data by focusing on incorporating domain-specific knowledge into RMs through model merging.
@@ 3.1 Reward Modeling
$\mathcal{L}_{\text{RM}}=-\log{[\sigma(r(x,y_{c}))-\sigma(r(x,y_{r}))]}$ (1)
@@
Model Reward Bench Auto-J Eval Best-of-16
Chat Chat-Hard Safety Reasoning Code Math Others GSM8K MBPP
@@
(a) LLaMA-2 RM 95.8 47.6 44.6 78.9 68.2 76.2 84.4 79.2 35.3 17.2
(b) FT on Auto-J Math 94.7 48.5 44.4 79.1 68.7 76.2 90.2 79.2 35.2 -
(c) FT on Auto-J Code 94.7 48.2 44.3 78.8 66.9 89.3 84.4 79.4 - 17.2
(d) Ours (+ MetaMath) 95.8 44.5 43.5 85.7 79.6 79.8 87.5 79.3 40.7 -
(e) Ours (+ MAmmoTH) 96.1 44.7 43.8 84.1 85.2 79.8 87.5 79.7 40.5 -
(f) Ours (+ Code Model) 96.1 45.6 43.9 84.3 71.8 82.1 87.5 79.7 - 17.2
Table 1: Performance comparison across various benchmarks. Row (a) represents our base LLaMA-2 7B (Touvron et al., 2023) reward model. Rows (b) and (c) show results after fine-tuning the LLaMA-2 RM using the test data from Auto-J Eval (Li et al., 2024) Math and Code subsets, respectively. We use $\dagger$ to denote training accuracy, as these values are derived from benchmark testing data used during training. Rows (d) to (f) demonstrate the performance of LLaMA-2 RM when merged with MetaMath-7B (Yu et al., 2024), MAmmoTH-7B (Yue et al., 2024a), and the Code Model, each with a weight factor of $\lambda=0.35$.
@@ 3.2 Model Merging
$\theta^{\text{SFT}}=\{\theta^{\text{SFT}}_{\text{emb}},\theta^{\text{SFT}}_{\text{trans}},\theta^{\text{SFT}}_{\text{dec}}\}$ (2)
@@
$\theta^{\text{RM}}=\{\theta^{\text{RM}}_{\text{emb}},\theta^{\text{RM}}_{\text{trans}},\theta^{\text{RM}}_{\text{reg}}\}$ (3)
@@
where $t_{i}$ is a common token to both models, $\theta_{\text{emb},t_{i}}$ is the corresponding embedding, and $\lambda$ is a hyperparameter controlling the weight of the SFT parameters, ranging from 0 to 1.
@@
$\theta^{\text{MERGE}}_{\text{emb},t_{i}}=\begin{cases}\theta^{\text{SFT}}_{\text{emb},t_{i}}&\text{If $t_{i}$ is unique to SFT}\\ \theta^{\text{RM}}_{\text{emb},t_{i}}&\text{If $t_{i}$ is unique to RM}\end{cases}$ (5)
@@ 4.1 Experimental Setup
To fine-tune the backbone of our reward model, we utilize the 10k SFT split from Alpacafarm (Dubois et al., 2023). For reward modeling, we employ the UltraFeedback (Cui et al., 2024). The details of training and these datasets are presented in Appendix A and D, respectively.
@@
For the math LLMs, we adopt the open-source models, MetaMath-7B (Yu et al., 2024) and MAmmoTH-7B (Yue et al., 2024a), both of which are fine-tuned from LLaMA-2-7B. For code generation LLM, since we could not find open-source models with detailed training information, we use OSS-Instruct and Magicoder-Evol-Instruct (Wei et al., 2024) to fine-tune LLaMA-2-7B ourselves. We refer to this model as the Code Model in the following sections. The details of code fine-tuning datasets and math models are presented in Appendix D, E, and F.
@@ 4.2 Evaluation
We evaluate the reward models using two benchmarks, RewardBench (Lambert et al., 2024) and Auto-J Eval (Li et al., 2024). These benchmarks provide paired instruction-completion data, with the preferred completion annotated as chosen and the other as rejected, using accuracy as the evaluation metric. We use the core set of RewardBench, focusing primarily on the reasoning category to evaluate the model’s abilities in code and mathematical reasoning. For Auto-J Eval, we use pair-wise testing data and categorize the dataset into three categories: code, math, and others, following Yuan et al. 2024. Additionally, to further test our reward model’s effectiveness in enhancing model performance through reranking, we conduct best-of-N sampling on zero-shot prompted responses from llama-2-7B-chat for GSM8K (Cobbe et al., 2021) and MBPP (Austin et al., 2021). None of the models used in the experiment were trained on these data sources. Details about the models and datasets are provided in Appendix D and E, while the hyperparameters used for best-of-N sampling are outlined in Appendix B.
Additionally, in DogeRM, determining an appropriate weight factor $\lambda$ depends on a small in-domain validation set. This raises an important question: Can fine-tuning the reward model on this small dataset match or even exceed the performance of our method in the target domain? To investigate this, we performed continuous fine-tuning of our LLaMA-2 RM using test data from Auto-J Eval and evaluated the newly fine-tuned RM on the remaining benchmarks to assess its effectiveness.
Lastly, We present results using a weight factor of $\lambda=0.35$ for the main findings, with an analysis of the impact of $\lambda$ detailed in section 4.4 and results for different values of $\lambda$ presented in Appendix H.
@@ 4.3 Results
The main results of RM benchmarks are shown in Table 1. Merging the LLaMA-2 RM with MetaMath-7B (Row (d)) and MAmmoTH-7B (Row (e)) improves math performance on RewardBench by 11.4% and 17%, respectively, and coding performance by 5.2% and 5.8%, respectively. Similar enhancements are seen on Auto-J Eval, with gains in both math and coding. Merging our LLaMA-2 RM with the Code Model (Row (f)) further improves coding performance on RewardBench and Auto-J Eval by 5.4% and 6%, respectively, along with noticeable improvements in math performance on both benchmarks. Although DogeRM enhances performance in the reasoning domain, there is no significant degradation in other domains. The specific role of domain knowledge is evident, as merging with the math model leads to greater improvements in the math domain than merging with the Code model, and vice versa.
@@
Figure 2 and Table 1 show the results, with accuracy improvements on GSM8K. At the best-of-16 setting, merging with math models (Rows (d) and (e)) improves GSM8K by 5%, while merging with the Code Model (Row (f)) maintains performance on MBPP without degradation. We attribute the modest improvement on MBPP to the low upper bound of reranking performance (indicated by the black line in Figure 2(b)), which constrains the potential gains from reranking in this task.
@@
Fine-tuning on Small Validation Dataset
Rows (b) and (c) of Table 1 show the results of fine-tuning our LLaMA-2 RM on the Auto-J Eval Math and Code test subsets, respectively. While fine-tuning improved performance on Auto-J Eval, it did not generalize well to other benchmarks. In contrast, using these datasets as a validation set to determine an appropriate $\lambda$ for merging resulted in better overall performance. For a detailed analysis of $\lambda$’s impact across different benchmarks, see Appendix H.
@@ 4.4 Analysis
To test the adaptability of DogeRM to different model architectures, we use an open-source Mistral-based (Jiang et al., 2023) RM (Ray2333, 2024) merging with Misrtral-based MAmmoTH2-7B-Plus (Yue et al., 2024b). Details of these models are presented in Appendix E. The results for $\lambda=0.35$ in reasoning domains on RM benchmarks and best-of-N sampling on GSM8K, with N=16, are presented in Table 2. Our method improves math performance by 30% on RewardBench and 3% on Auto-J Eval. Additionally, we enhance reranking performance on GSM8K by 2.4%. These results demonstrate the adaptability of our methods to different model architectures. The results for different $\lambda$ are presented in Appendix H.
@@
Integrating Multiple Domains
To evaluate DogeRM’s capability of integrating knowledge from multiple domains, we experimented by merging MAmmoTH (Yue et al., 2024a) and the Code model into LLaMA-2 RM. We heuristically set the weight factors for MAmmoTH, the Code model, and LLaMA-2 RM at 0.2, 0.2, and 0.6, respectively. The evaluation results, presented in Table 3, indicate that merging models from multiple related domains can indeed enhance performance in those domains.
Model Reward Bench Auto-J Eval Best-of-16
Code Math Code Math GSM8K MBPP
LLaMA-2 RM 78.9 68.2 76.2 84.2 35.3 17.2
+ Math & Code 83.0 85.2 81.0 87.5 39.5 17.0
Table 3: Performance of merging LLaMA-2 RM with MAmmoTH-7B and the Code model on various benchmarks and best-of-16 results. Our methods show improvements across RM benchmarks and in best-of-16 sampling on GSM8K.
@@ 5 Conclusion
In this work, we introduce a novel approach, DogeRM, which integrates domain knowledge into RM by merging it with the domain-specific SFT models. We demonstrate that DogeRM enhances performance on math and coding benchmarks and can be generalized to different model architectures. A series of analyses show that DogeRM effectively affects the reward signal corresponding to chosen and rejected prompts. The results highlight DogeRM’s potential to enhance model alignment and generation verification through model merging, offering promising results across various benchmarks.
@@
There are several limitations in our work: (1) Our framework has been tested exclusively in the math and coding domains, leaving other areas such as medicine, finance, and law unexplored. In Section 4.4, we demonstrated the effectiveness of merging domain-specific models for math and coding into RM. However, the integration of models from multiple orthogonal domains remains an area for future investigation. (2) Our method was tested exclusively on 7B models, and we have not evaluated its performance on models of larger or smaller sizes. (3) While our framework is compatible with various merging techniques, such as TIES-Merge (Yadav et al., 2024), we have not thoroughly examined the impact of these more advanced methods. In this work, we focused on demonstrating the core idea—improving RM performance in a target domain by merging with a domain-specific language model. To keep our approach straightforward, we used weighted averaging, which, in the case of two models, can be understood as a form of task arithmetic. Despite the simplicity of this method, it has already led to notable performance gains. However, the effectiveness of more sophisticated merging techniques remains unexplored, and we leave this investigation for future work. (4) The models being merged must share the same architecture, a limitation common to most model merging algorithms (Wortsman et al., 2022; Ilharco et al., 2023; Yadav et al., 2024). Recently, evolutionary model merging (Akiba et al., 2024) has been proposed as a solution for merging models with different architectures. Investigating the merging of models with varying architectures remains a topic for future research. (5) Due to the sensitivity of RLHF to hyperparameter choices and our limited computational resources, we did not implement RLHF algorithms such as PPO (Schulman et al., 2017) or RLOO (Ahmadian et al., 2024) in this work. Exploring the integration of DogeRM within RLHF frameworks is left for future work.